# Data-Driven Reduced-Order Models for Port-Hamiltonian Systems with Operator Inference

### v1

@@ 1 Introduction
For systems with certain structure, such as Lagrangian, Hamiltonian and pH systems, failure to satisfy the structure at the ROM level can result in unphysical solutions and unstable behavior [34, 12]. For instance, the reduced-order operators for ${\mathbf{J}}$ and ${\mathbf{R}}$ in (1) may lose their respective skew-symmetry or symmetric semi-positive definiteness properties. To address this issue, structure-preserving MOR techniques have been introduced. Similar to other types of MOR, structure-preserving ROMs can be constructed either intrusively or non-intrusively. Intrusive methods include the POD-based Galerkin projection method for Hamiltonian systems [22, 35, 6], POD/$\mathcal{H}_{2}$-based Petrov-Galerkin projection method for pH systems [8, 13], proper symplectic decomposition (PSD) for canonical Hamiltonian systems [40], and RB-ROMs for Hamiltonian systems [1, 27, 37, 28]. Additionally, to efficiently evaluate nonlinear terms in ROMs, the discrete empirical interpolation method (DEIM), originally developed in [7, 14] for interpolating the nonlinear terms over selected sparse sample points, has been extended in [38] to the structure-preserving setting for Hamiltonian ROMs. Unlike classical model reduction approaches that use linear subspaces, the authors in [44, 48, 11] combine quadratic or nonlinear manifold learning with symplectic auto-encoders to construct an intrusive ROM for Hamiltonian systems. Non-intrusive structure-preserving MOR techniques, on the other hand, construct ROMs by using only snapshot data and some prior knowledge of the dynamical system. Hence they are sometimes referred to as ‘glass box’ approaches. Recent work includes learning of canonical Hamiltonian ROMs based on operator inference, where the linear part of the Hamiltonian gradient is inferred through constrained least-squares solutions [45]. Non-canonical Hamiltonian operator inference ROMs have been developed in [23], assuming the entire Hamiltonian function is known and inferring the reduced operator associated with the linear differential operator. Gradient-preserving operator inference (GP-OpInf) has been developed for conservative or dissipative systems in [21].
@@
In this paper, we propose a pH operator inference (pH-OpInf) method to construct ROMs for both linear and nonlinear pH systems, using snapshot data of the state, external input, and output, along with knowledge of the functional form of the Hamiltonian. The main contributions of our work include: (i) we propose two optimization formulations to learn the structure-preserving reduced-order operators for constructing ROMs of pH systems. In particular, the optimization in the first formulation is decoupled into two subtasks in the second formulation to facilitate its solution; (ii) we incorporate a hyper-reduction method into the structure-preserving ROMs for nonlinear pH systems; (iii) we derive a priori error estimates for the reduced-order approximations for both state and output variables.
The remainder of this paper is organized as follows. In Section 2, we review a Galerkin-projection-based ROM as the state-of-the-art intrusive MOR method for reducing the pH system (1), and employ DEIM to ensure efficient simulation of the nonlinear ROM. Section 3 proposes the new pH-OpInf method and suggests two optimization problems for finding the reduced-order operators. In Section 4, we analyze the resulting structure-preserving ROMs and derive a priori error estimates for both state and output approximations. The effectiveness of the proposed ROMs is numerically demonstrated through two examples in Section 5. Finally, some concluding remarks are drawn in Section 6.
@@ 2 Background on Projection-Based Reduced-Order Models for the Port-Hamiltonian System
One widely used nonlinear MOR technique is POD [10], which seeks a set of reduced basis vectors from snapshot data. In practice, the basis vectors are often computed using singular value decomposition (SVD) on the array of snapshots, and the associated left singular vectors are truncated to form the reduced basis. The truncation error provides empirical guidance on choosing the number of reduced basis vectors. By approximating the state variables in the subspace spanned by the reduced basis and projecting the original system onto either the same subspace or another low-dimensional space, one can construct a Galerkin or Petrov-Galerkin projection-based ROM. This workflow is intrusive, as it requires access to the matrices and vectors of the full-order model (FOM) to build the ROM. When applied to Hamiltonian or pH systems, such ROMs may alter the structural properties of the FOM and result in inaccurate and unphysical solutions. This limitation motivates the development of structure-preserving ROMs, see examples in [24, 13, 22]. In the following, we focus on Galerkin projection and the design of a structure-preserving ROM for the pH system (1).
@@ 2.1 Intrusive Structure-Preserving ROMs
where ${\mathbf{x}}(t^{i})$ is the state vector at the discrete time instances $t^{i}$, for $i=0,1,\dots,s$. For simplicity, we assume the snapshots are uniformly distributed, that is, $t^{i}=i\Delta t$ with a constant $\Delta t$ that could be the time step size used in full-order simulations or a checkpoint at which solutions are stored. The POD basis matrix is denoted as ${\boldsymbol{\Phi}}\coloneqq[\phi_{1},\phi_{2},\dots,\phi_{r}]$, where $r\leq n$ is the dimension of the POD basis and $\phi_{i}$ is the $i$th left singular vector of $\mathbf{X}$. Note that ${\boldsymbol{\Phi}}^{\intercal}{\boldsymbol{\Phi}}={\bf I}_{r}$. The reduced approximation employs the ansatz $\hat{{\mathbf{x}}}(t)={\boldsymbol{\Phi}}{\mathbf{x}}_{r}(t)$, where ${\mathbf{x}}_{r}(t)\in\mathbb{R}^{r}$ is the unknown POD coefficient vector.
@@
Replacing ${\mathbf{J}}$, ${\mathbf{R}}$ and ${\mathbf{B}}$ in (4) with $\widetilde{{\mathbf{J}}}_{r}$, $\widetilde{{\mathbf{R}}}_{r}$ and $\widetilde{{\mathbf{B}}}_{r}={\boldsymbol{\Phi}}^{\intercal}{\mathbf{B}}$, respectively, and noticing that
@@
Because the dimension $r$ of the ROM is typically much smaller than $n$, simulating the SP-G-ROM is generally more computationally efficient than simulating the FOM. However, when $\nabla_{{\mathbf{x}}_{r}}H_{r}({\mathbf{x}}_{r}(t))$ is nonlinear, the overall computational complexity still depends on $n$. Therefore, an additional approximation, termed hyper-reduction, is needed to accelerate the ROM simulation.
@@ 2.2 Hyper-Reduction of Nonlinear Hamiltonian ROMs
The calculation of $\nabla_{{\mathbf{x}}_{r}}N_{r}({\mathbf{x}}_{r}(t))$ is expensive for large $n$. To reduce the computational cost, we apply the approach developed in [38], outlined below, that extends DEIM to such a nonlinear Hamiltonian function. First, without loss of generality, $N_{r}({\mathbf{x}}_{r}(t))$ can be written as
@@
$M_{J}=[J_{h}({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}(t^{0})){\boldsymbol{\Phi}},\dots,J_{h}({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}(t^{s})){\boldsymbol{\Phi}}]\in\mathbb{R}^{d\times r(s+1)}.$
The standard DEIM as in [14, 19] is then applied to extract the nonlinear basis ${\boldsymbol{\Psi}}\in\mathbb{R}^{d\times m}$ from $M_{J}$ and determine an interpolation matrix ${\mathbf{P}}:=[{\mathbf{e}}_{\wp_{1}},\dots,{\mathbf{e}}_{\wp_{m}}]$, where ${\mathbf{e}}_{i}$ is the $i$th unit vector of $\mathbb{R}^{d}$ and $\{\wp_{1},\dots,\wp_{m}\}\subset\{1,\dots,d\}$ are the interpolation indices. Letting $\mathbb{P}={\boldsymbol{\Psi}}({\mathbf{P}}^{\intercal}{\boldsymbol{\Psi}})^{-1}{\mathbf{P}}^{\intercal}$, the nonlinear term $N_{r}({\mathbf{x}}_{r}(t))$ is approximated by
@@
and we have the same initial condition as in (7). Calculating the rate of change in $H_{h_{r}}({\mathbf{x}}_{r}(t))$ reveals that the dissipation inequality is satisfied at the reduced-order level. Since $\nabla_{{\mathbf{x}}_{r}}H_{h_{r}}({\mathbf{x}}_{r}(t))$ is calculated with the DEIM interpolation, the ROM’s online simulation is faster than the SP-G-ROM. For a detailed computational complexity analysis, see [38].
@@ 3.2 Two Optimization Problems for pH-OpInf
Based on these generated data from (11) and (12), one can infer ${\mathbf{J}}_{r}$, ${\mathbf{R}}_{r}$ and ${\mathbf{B}}_{r}$ for the pH-OpInf-ROM in (10) through optimization. An iterative algorithm proposed in [36] recasts the entire coefficient matrix $\left[\begin{smallmatrix}{\mathbf{J}}_{r}-{\mathbf{R}}_{r}&{\mathbf{B}}_{r}\\ -{\mathbf{B}}_{r}^{\intercal}&\mathbf{0}\end{smallmatrix}\right]$ into $\boldmath\widetilde{J}-\widetilde{R}$, and then learns the skew-symmetric matrix $\boldmath\widetilde{J}$ and the symmetric and positive semi-definite matrix $\boldmath\widetilde{R}$ iteratively. However, using the learned $\boldmath\widetilde{J}$ and $\boldmath\widetilde{R}$ can not ensure that the (2,2)-block in the coefficient matrix is zero with machine precision. Moreover, the iterative algorithm requires a delicate initial guess to ensure convergence. In this work, we identify the reduced operators with appropriate structural properties through a subsystem-based optimization rather than a monolithic approach. Next, we propose and analyze two approaches.
@@
First, we define
${\mathbf{D}}_{r}\coloneqq{\mathbf{J}}_{r}-{\mathbf{R}}_{r},\quad\text{where}\quad{\mathbf{J}}_{r}\coloneqq\frac{1}{2}({\mathbf{D}}_{r}-{\mathbf{D}}_{r}^{\intercal})\quad\text{and}\quad{\mathbf{R}}_{r}\coloneqq-\frac{1}{2}({\mathbf{D}}_{r}+{\mathbf{D}}_{r}^{\intercal}).$
Then we find ${\mathbf{D}}_{r}$ and ${\mathbf{B}}_{r}$ simultaneously by minimizing an objective function that incorporates the residuals of both the state equation (10a) and the output equation (10b) at the reduced level. The residuals are balanced using a weighting parameter $\lambda_{w}>0$ and we get
@@
This optimization has a semi-definite constraint and can be solved by an interior point method [26].
@@ 3.3 Data-driven pH-OpInf-ROMs
In both optimization problems, since $\frac{1}{2}({\mathbf{D}}_{r}+{\mathbf{D}}_{r}^{\intercal})\preccurlyeq\mathbf{0}$, the inferred matrix ${\mathbf{R}}_{r}$ must be semi-positive definite; and ${\mathbf{R}}_{r}$ and ${\mathbf{J}}_{r}$ are symmetric and skew-symmetric, respectively, due to their definitions. After substituting them into the ROM (10), we obtain the pH-OpInf-ROMs that are structure-preserving. According to the optimization formulations, we name them respectively by pH-OpInf-W and pH-OpInf-R.
Correspondingly, after incorporating the DEIM discussed in Section 2.2, we obtain the pH-OpInf-DEIM model:
@@ 4 Error Estimates for Port-Hamiltonian Operator Inference ROMs
Four terms appear in the error bound: the first term $\int_{0}^{T}\|{\mathbf{x}}-{\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}\|^{2}\,{\rm d}t$ measures the projection error caused by projecting ${\mathbf{x}}(t)$ onto the subspace spanned by the POD basis ${\boldsymbol{\Phi}}$, which has been analyzed thoroughly in [32, 46]; the second one $\int_{0}^{T}\|\accentset{\mbox{\large.}}{{\mathbf{x}}}-\mathcal{D}_{t}[{\mathbf{x}}]\|^{2}\,{\rm d}t$ is the data error caused by generating the time derivative snapshots using a finite difference scheme; the third one $\int_{0}^{T}\|{\boldsymbol{\Phi}}^{\intercal}\mathcal{D}_{t}[{\mathbf{x}}]-{\mathbf{D}}_{r}{\boldsymbol{\Phi}}^{\intercal}\nabla_{{\mathbf{x}}}H({\mathbf{x}})-{\mathbf{B}}_{r}{\mathbf{u}}(t)\|^{2}\,{\rm d}t$ represents the optimization error due to the imperfect learning of ${\mathbf{D}}_{r}$ and ${\mathbf{B}}_{r}$ from data; and the forth one ${\int_{0}^{T}\|(\mathbf{I}-\mathbb{P})J_{h}({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}){\boldsymbol{\Phi}}\|^{2}\,{\rm d}t}$ is the hyper-reduction error caused by the DEIM approximation of the nonlinear Jacobian, which is used to accelerate the evaluation of the gradient of reduced Hamiltonian (9). Next, we present the proof.
@@
For the last term, following [15, Thm. 3.1] and [38, Thm. 3.3], its linear part can be bounded by

### v2

@@ 1 Introduction
For systems with certain structure, such as Lagrangian, Hamiltonian and pH systems, failure to satisfy the structure at the ROM level can result in unphysical solutions and unstable behavior [14, 15]. For instance, the reduced-order operators for ${\mathbf{J}}$ and ${\mathbf{R}}$ in (1) may lose their respective skew-symmetry or symmetric semi-positive definiteness properties. To address this issue, structure-preserving MOR techniques have been introduced. Similar to other types of MOR, structure-preserving ROMs can be constructed either intrusively or non-intrusively. Intrusive methods include the POD-based Galerkin projection method for Hamiltonian systems [16, 17, 18], POD/$\mathcal{H}_{2}$-based Petrov-Galerkin projection method for pH systems [19, 20], proper symplectic decomposition (PSD) [21], a variationally consistent approach for canonical Hamiltonian systems [22], and RB-ROMs for Hamiltonian systems [23, 24, 25, 26]. Additionally, to efficiently evaluate nonlinear terms in ROMs, the discrete empirical interpolation method (DEIM), originally developed in [27, 28] for interpolating the nonlinear terms over selected sparse sample points, has been extended in [29] to the structure-preserving setting for Hamiltonian ROMs. Unlike classical model reduction approaches that use linear subspaces, the authors in [30, 31, 32] combine quadratic or nonlinear manifold learning with symplectic auto-encoders to construct an intrusive ROM for Hamiltonian systems. Non-intrusive structure-preserving MOR techniques, on the other hand, construct ROMs by using only snapshot data and some prior knowledge of the dynamical system. Hence, they are sometimes referred to as ‘glass box’ approaches. Recent work includes learning of canonical Hamiltonian ROMs based on operator inference, where the linear part of the Hamiltonian gradient is inferred through constrained least-squares solutions [33]. Non-canonical Hamiltonian operator inference ROMs have been developed in [34], assuming the entire Hamiltonian function is known and inferring the reduced operator associated with the linear differential operator. Gradient-preserving operator inference (GP-OpInf) has been developed for conservative or dissipative systems in [35]. A tensor-based operator inference approach has been introduced in [36] to construct a ROM with the Hamiltonian structure while capturing parametric dependencies.
@@
In this paper, we propose a pH operator inference (pH-OpInf) method to construct ROMs for linear and nonlinear pH systems, using snapshot data of the state, external input, and output, along with knowledge of the functional form of the Hamiltonian. The main contributions of our work include:
  1. (i)
    We propose two optimization formulations to learn the structure-preserving reduced-order operators for constructing ROMs of pH systems. In particular, the optimization in the first formulation is decoupled into two subtasks in the second formulation to facilitate its solution.
  2. (ii)
    We incorporate hyper-reduction into the structure-preserving ROMs for nonlinear pH systems.
  3. (iii)
    We derive a priori error estimates for the reduced-order approximations of the inferred structure-preserving ROM for both state and output variables.
The remainder of this paper is organized as follows. In Section 2, we review a Galerkin projection-based ROM as the state-of-the-art intrusive MOR method for reducing the pH system (1), and employ DEIM to ensure efficient simulation of the nonlinear ROM. Section 3 proposes the new pH-OpInf method and suggests two optimization problems to find the reduced-order operators. In Section 4, we analyze the resulting structure-preserving ROMs and derive a priori error estimates for both state and output approximations. The effectiveness of the proposed ROMs is numerically demonstrated through two examples in Section 5. Finally, some concluding remarks are drawn in Section 6.
@@ 2 Background on Projection-Based Reduced-Order Models for the Port-Hamiltonian System
The POD method [39] has been widely used in MOR to find a set of reduced basis vectors from snapshot data, which are then used to construct the ROM evolution equations. In practice, the basis vectors are often computed using singular value decomposition (SVD) on the array of snapshots, and the associated left singular vectors are truncated to form the reduced basis. The truncation error provides empirical guidance on choosing the number of reduced basis vectors. By approximating the state variables in the subspace spanned by the reduced basis and projecting the original system onto either the same subspace or another low-dimensional space, one can construct a Galerkin or Petrov-Galerkin projection-based ROM. This workflow is intrusive, as it requires access to the matrices and vectors of the full-order model (FOM) to build the ROM. When applied to Hamiltonian or pH systems, such ROMs may alter the structural properties of the FOM and result in inaccurate and unphysical solutions. This limitation motivates the development of structure-preserving ROMs, see examples in [40, 20, 16]. In the following, we focus on Galerkin projection and the design of a structure-preserving ROM for the pH system (1).
@@ 2.1 Intrusive Structure-Preserving ROMs
where ${\mathbf{x}}(t^{i})$ is the state vector at the discrete time instances $t^{i}$, for $i=0,1,\dots,s$. For simplicity, we assume the snapshots are uniformly distributed, that is, $t^{i}=i\Delta t$ with a constant $\Delta t$ that could be the time step size used in full-order simulations or a checkpoint at which solutions are stored. The POD basis matrix is denoted as ${\boldsymbol{\Phi}}\coloneqq[\phi_{1},\phi_{2},\dots,\phi_{r}]$, where $\phi_{i}$ is the $i$th left singular vector of $\mathbf{X}$ associated with the singular value $\sigma_{i}$, and $r$ is the dimension of the POD basis that is less than the rank $d$ of ${\mathbf{X}}$. The dimension $r$ is typically selected to ensure that the POD basis captures a significant portion of the snapshot energy. In particular, given a prescribed tolerance $\tau$, $r$ is chosen such that $\frac{\sum_{i=1}^{r}\sigma_{i}^{2}}{\sum_{i=1}^{d}\sigma_{i}^{2}}\geq\tau$. Note that ${\boldsymbol{\Phi}}^{\intercal}{\boldsymbol{\Phi}}={\bf I}_{r}$. The reduced approximation employs the ansatz $\hat{{\mathbf{x}}}(t)={\boldsymbol{\Phi}}{\mathbf{x}}_{r}(t)$, where ${\mathbf{x}}_{r}(t)\in\mathbb{R}^{r}$ is the unknown POD coefficient vector.
@@
Replacing ${\boldsymbol{\Phi}}^{\intercal}{\mathbf{J}}$ and ${\boldsymbol{\Phi}}^{\intercal}{\mathbf{R}}$ in (4) with $\widetilde{{\mathbf{J}}}_{r}{\boldsymbol{\Phi}}^{\intercal}$ and $\widetilde{{\mathbf{R}}}_{r}{\boldsymbol{\Phi}}^{\intercal}$, respectively, defining $\widetilde{{\mathbf{B}}}_{r}={\boldsymbol{\Phi}}^{\intercal}{\mathbf{B}}$, and noticing that
@@
Since the dimension $r$ of the ROM is typically much smaller than $n$, simulating the SP-G-ROM is generally more computationally efficient than simulating the FOM. However, when $\nabla_{{\mathbf{x}}_{r}}H_{r}({\mathbf{x}}_{r}(t))$ is nonlinear, the overall computational complexity still depends on $n$. Therefore, an additional approximation, termed hyper-reduction, is needed to accelerate the ROM simulation.
@@ 2.2 Hyper-Reduction of Nonlinear Hamiltonian ROMs
The calculation of $\nabla_{{\mathbf{x}}_{r}}N_{r}({\mathbf{x}}_{r}(t))$ is expensive for large $n$. To reduce the computational cost, we apply the approach developed in [29], outlined below, that extends DEIM to such a nonlinear Hamiltonian function. Note that both standard DEIM [28, 41] and Q-DEIM [41] can be used to select interpolation points. We choose the latter because of its enhanced stability and accuracy. Since Q-DEIM is a variant of DEIM, we will continue to use ‘DEIM’ to refer to the hyper-reduction approach and label the associated hyper-reduced ROMs.
First, without loss of generality, $N_{r}({\mathbf{x}}_{r}(t))$ can be written as
@@
$M_{J}=[J_{h}({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}(t^{0})){\boldsymbol{\Phi}},\dots,J_{h}({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}(t^{s})){\boldsymbol{\Phi}}]\in\mathbb{R}^{d\times r(s+1)},$
which is used to determine the nonlinear basis ${\boldsymbol{\Psi}}\in\mathbb{R}^{d\times m}$ and the interpolation matrix ${\mathbf{P}}:=[{\mathbf{e}}_{\wp_{1}},\dots,{\mathbf{e}}_{\wp_{m}}]$ using DEIM, where ${\mathbf{e}}_{i}$ is the $i$th unit vector of $\mathbb{R}^{d}$ and $\{\wp_{1},\dots,\wp_{m}\}\subset\{1,\dots,d\}$ are the interpolation indices. Letting $\mathbb{P}={\boldsymbol{\Psi}}({\mathbf{P}}^{\intercal}{\boldsymbol{\Psi}})^{-1}{\mathbf{P}}^{\intercal}$, the nonlinear term $N_{r}({\mathbf{x}}_{r}(t))$ is approximated by
@@
and we have the same initial condition as in (7). Calculating the rate of change in $H_{h_{r}}({\mathbf{x}}_{r}(t))$ reveals that the dissipation inequality is satisfied at the reduced-order level. Since $\nabla_{{\mathbf{x}}_{r}}H_{h_{r}}({\mathbf{x}}_{r}(t))$ is calculated with the help of interpolation, the ROM’s online simulation is faster than the SP-G-ROM. For a detailed computational complexity analysis, see [29].
@@ 3.2 Two Optimization Problems for pH-OpInf
Based on these data generated from (11) and (12), one can infer ${\mathbf{J}}_{r}$, ${\mathbf{R}}_{r}$ and ${\mathbf{B}}_{r}$ for the pH-OpInf-ROM in (10) through optimization. An iterative algorithm proposed in [38] recasts the entire coefficient matrix $\left[\begin{smallmatrix}{\mathbf{J}}_{r}-{\mathbf{R}}_{r}&{\mathbf{B}}_{r}\\ -{\mathbf{B}}_{r}^{\intercal}&\mathbf{0}\end{smallmatrix}\right]$ into $\boldmath\widetilde{J}-\widetilde{R}$, and then learns the skew-symmetric matrix $\boldmath\widetilde{J}$ and the symmetric and positive semi-definite matrix $\boldmath\widetilde{R}$ iteratively. However, using the learned $\boldmath\widetilde{J}$ and $\boldmath\widetilde{R}$ cannot ensure that the (2,2) block in the coefficient matrix is zero with machine precision. Moreover, the iterative algorithm requires a delicate initial guess to ensure convergence. In this work, we identify the reduced operators with appropriate structural properties through a subsystem-based optimization rather than a monolithic approach.
Next, we propose and analyze two approaches. In both, we first determine ${\mathbf{D}}_{r}$, then post process it to obtain ${\mathbf{J}}_{r}\coloneqq\frac{1}{2}({\mathbf{D}}_{r}-{\mathbf{D}}_{r}^{\intercal})$ and ${\mathbf{R}}_{r}\coloneqq-\frac{1}{2}({\mathbf{D}}_{r}+{\mathbf{D}}_{r}^{\intercal})$ so that we again have ${\mathbf{D}}_{r}={\mathbf{J}}_{r}-{\mathbf{R}}_{r}$.
@@
We find ${\mathbf{D}}_{r}$ and ${\mathbf{B}}_{r}$ simultaneously by minimizing an objective function that incorporates the residuals of the state equation (10a) and the output equation (10b) at the reduced level. The residuals are balanced using a weighting parameter $\lambda_{w}>0$ and we get
@@
This optimization has a semi-definite constraint and can be solved by an interior point method [42]. Operator inference implementations often use regularization terms to solve unconstrained optimization problems. However, we did not observe any additional numerical improvements for our constrained optimization in the test cases investigated in Section 5, so we did not implement regularization.
@@ 3.3 Data-driven pH-OpInf-ROMs
In both optimization problems, since $\frac{1}{2}({\mathbf{D}}_{r}+{\mathbf{D}}_{r}^{\intercal})\preccurlyeq\mathbf{0}$, the inferred matrix ${\mathbf{R}}_{r}$ must be positive semi-definite; and ${\mathbf{R}}_{r}$ and ${\mathbf{J}}_{r}$ are symmetric and skew symmetric, respectively, due to their definitions. After substituting them into the ROM (10), we obtain the pH-OpInf-ROMs that are structure-preserving. Following the terminology of the optimization formulations, we refer to them as pH-OpInf-W and pH-OpInf-R.
Correspondingly, after incorporating the hyper-reduction approach discussed in Section 2.2, we obtain the pH-OpInf-DEIM model:
@@ 4 Error Estimates for Port-Hamiltonian Operator Inference ROMs
Four terms appear in the error bound: the first term $\int_{0}^{T}\|{\mathbf{x}}-{\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}\|^{2}\,{\rm d}t$ measures the projection error caused by projecting ${\mathbf{x}}(t)$ onto the subspace spanned by the POD basis ${\boldsymbol{\Phi}}$, which has been analyzed thoroughly in [45, 46]; the second one $\int_{0}^{T}\|\accentset{\mbox{\large.}}{{\mathbf{x}}}-\mathcal{D}_{t}[{\mathbf{x}}]\|^{2}\,{\rm d}t$ is the data error caused by generating the time derivative snapshots using a finite difference scheme; the third one $\int_{0}^{T}\|{\boldsymbol{\Phi}}^{\intercal}\mathcal{D}_{t}[{\mathbf{x}}]-{\mathbf{D}}_{r}{\boldsymbol{\Phi}}^{\intercal}\nabla_{{\mathbf{x}}}H({\mathbf{x}})-{\mathbf{B}}_{r}{\mathbf{u}}(t)\|^{2}\,{\rm d}t$ represents the optimization error, which comprises the model error due to fitting the reduced-order operators ${\mathbf{D}}_{r}$ and ${\mathbf{B}}_{r}$ using the projected FOM data, and the numerical error arising from solving the optimization problem by numerical algorithms; and the forth one ${\int_{0}^{T}\|(\mathbf{I}-\mathbb{P})J_{h}({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}{\mathbf{x}}){\boldsymbol{\Phi}}\|^{2}\,{\rm d}t}$ is the hyper-reduction error caused by the DEIM approximation of the nonlinear Jacobian, which is used to accelerate the evaluation of the gradient of reduced Hamiltonian (9). Next, we present the proof.
@@
For the last term, following [43, Thm. 3.1] and [29, Thm. 3.3], its linear part can be bounded by
@@
Remark 1.
The re-projection method has been introduced in [47] to reduce optimization errors for operator inference. When the FOM drift terms are polynomial and the time derivative data is discretized in the same manner as the FOM, the re-projection approach ensures that the inferred ROM recovers the intrusive Galerkin-projection ROM and the optimization error is zero. However, this does not hold for port-Hamiltonian systems. Suppose the re-projection method generates a sequence of approximate state vectors $\{\widetilde{{\mathbf{x}}}_{0},\widetilde{{\mathbf{x}}}_{1},\ldots,\widetilde{{\mathbf{x}}}_{s}\}$, in which $\widetilde{{\mathbf{x}}}_{n+1}$ is obtained from (1a) after approximating the time derivative by $\mathcal{D}_{t}$ and taking $\widetilde{{\mathbf{x}}}_{n}$ as the initial data. Then it satisfies
$\mathcal{D}_{t}[{\boldsymbol{\Phi}}^{\intercal}\,\widetilde{{\mathbf{x}}}_{n}]={\boldsymbol{\Phi}}^{\intercal}{\mathbf{D}}\nabla_{{\mathbf{x}}}H({\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}\,\widetilde{{\mathbf{x}}}_{n})+{\boldsymbol{\Phi}}^{\intercal}{\mathbf{B}}{\mathbf{u}}_{n}.$ (30)
Define $\widetilde{\mathbf{X}}\coloneqq{\boldsymbol{\Phi}}{\boldsymbol{\Phi}}^{\intercal}\left[\widetilde{{\mathbf{x}}}_{0},\widetilde{{\mathbf{x}}}_{1},\ldots,\widetilde{{\mathbf{x}}}_{s}\right]\in\mathbb{R}^{r\times(s+1)}$. Assuming that ${\mathbf{B}}_{r}={\boldsymbol{\Phi}}^{\intercal}{\mathbf{B}}$ has been obtained, we next find ${\mathbf{D}}_{r}\in\mathbb{R}^{r\times r}$ by minimizing
$\frac{1}{2}\left\|{\boldsymbol{\Phi}}^{\intercal}\,\mathcal{D}_{t}\left[\widetilde{{\mathbf{X}}}\right]-{\mathbf{D}}_{r}{\boldsymbol{\Phi}}^{\intercal}\nabla_{{\mathbf{x}}}H\left(\widetilde{{\mathbf{X}}}\right)-{\mathbf{B}}_{r}{\mathbf{U}}\right\|_{F}^{2}\quad\text{s.t.}\quad\frac{1}{2}({\mathbf{D}}_{r}+{\mathbf{D}}_{r}^{\intercal})\preccurlyeq\mathbf{0}.$
Because of the relation in (30) and $\mathcal{D}_{t}[{\boldsymbol{\Phi}}^{\intercal}\,\widetilde{{\mathbf{X}}}]={\boldsymbol{\Phi}}^{\intercal}\,\mathcal{D}_{t}[\widetilde{{\mathbf{X}}}]$, it is equivalent to minimize
$\frac{1}{2}\left\|{\boldsymbol{\Phi}}^{\intercal}{\mathbf{D}}\nabla_{{\mathbf{x}}}H\left(\widetilde{{\mathbf{X}}}\right)-{\mathbf{D}}_{r}{\boldsymbol{\Phi}}^{\intercal}\nabla_{{\mathbf{x}}}H\left(\widetilde{{\mathbf{X}}}\right)\right\|_{F}^{2}\quad\text{s.t.}\quad\frac{1}{2}({\mathbf{D}}_{r}+{\mathbf{D}}_{r}^{\intercal})\preccurlyeq\mathbf{0}.$
In general, the minimizer ${\mathbf{D}}_{r}$ solved from the optimization can not reduce the objective function to zero.